OpenStudy (anonymous):

Two positive numbers have a difference of 4 and a product of 96. What are the numbers? 12 and 8 16 and 6 16 and 12 18 and 14

OpenStudy (anonymous):

(y+4)y = 96 y^2 + 4y - 96 = 0 (y-8)(y+12) = 0 Positive solution: y = 8 If y = 8 then 8x = 96 x = ?

OpenStudy (anonymous):

a